OpenStudy (anonymous):

solve the equation m/-6+6=41 m=???

OpenStudy (hba):

m=41

OpenStudy (anonymous):

\[\huge \frac{m}{-6}+6=41\] multiply it by -6 \[\huge m-36=-246\] \[\huge m=-246+36 \\ \huge m=-210\]